# Case2Code: Learning Inductive Reasoning with Synthetic Data

### v1

@@ Abstract
Complex reasoning is an impressive ability shown by large language models (LLMs). Most LLMs are skilled in deductive reasoning, such as chain-of-thought prompting or iterative tool-using to solve challenging tasks step-by-step. In this paper, we hope to focus on evaluating and teaching LLMs to conduct inductive reasoning, that is, LLMs are supposed to infer underlying rules by observing examples or sequential transformations. However, collecting large-scale and diverse human-generated inductive data is challenging. We focus on data synthesis in the code domain and propose a Case2Code task by exploiting the expressiveness and correctness of programs. Specifically, we collect a diverse set of executable programs, synthesize input-output transformations for each program, and force LLMs to infer the underlying code implementations based on the synthetic I/O cases. We first evaluate representative LLMs on the synthesized Case2Code task and demonstrate that the Case-to-code induction is challenging for LLMs. Then, we synthesize large-scale Case2Code training samples to train LLMs to perform inductive reasoning. Experimental results show that such induction training benefits not only in distribution Case2Code performance but also enhances various coding abilities of trained LLMs, demonstrating the great potential of learning inductive reasoning via synthetic data.111Code and datasets will be available at https://github.com/choosewhatulike/case2code.
Case2Code: Learning Inductive Reasoning with Synthetic Data
@@ 1 Introduction
The success of large language models (LLMs), exemplified by GPT-4 OpenAI (2023) has revolutionized the AI community. One of the most impressive abilities of LLMs is the deductive reasoning ability via chain-of-thoughts Wei et al. (2022b), exemplified by solving mathematical reasoning tasks such as GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021). The major contribution to the high-level performances of reasoning problem solving is to generate and train on deductive reasoning paths via chain-of-thoughts (CoTs) using LLMs Yu et al. (2023); Mitra et al. (2024). Through various searching strategies and prompting algorithms Chen et al. (2024); Wang et al. (2023a); Yao et al. (2023); Wang et al. (2022b), LLMs can synthesize high-quality CoTs and perform deduction reasoning in many domains.
Despite the success of deducing chain-of-thought reasoning, LLMs are rarely trained to perform inductive reasoning. Inductive reasoning is a fundamental cognitive process for humans, playing a crucial role in learning, problem-solving, and scientific discovery. For instance, famous scientists make inductive judgments from actual physical phenomena, such as Newton’s Law of Motion, and Kepler’s Law based on data from Tycho. It is essential to teach artificial intelligence systems utilizing inductive reasoning to find underlying rules from facts, transformations, and logs. In this paper, to equip LLMs with such inductive reasoning ability, we introduce Case2Code, a diverse and challenging inductive reasoning task for LLMs.
Specific and limited inductive reasoning has been studied in the machine learning field. Works such as DEER Yang et al. (2022) usually design a common knowledge induction process and challenge neural models to reason inductively to find the hidden fact. Works such as DreamCoder Ellis et al. (2021) and DeepCoder Balog et al. (2016), usually synthesize a small-scale toy task to train a domain-specific model, such as learning to summarize the list operations from given list change logs.
@@
Different from these inductive reasoning tasks, we focus on large-scale data synthesis with programs in the real world. In Case2Code, inductive reasoning samples are synthesized from real-world productive functions, which are closer to the actual distribution of general LLM applications and production. Specifically, the Case2Code challenge requires the LLM to infer the underlying program based on several input-to-output cases generated by the real-world program. In Case2Code learning, LLMs are supposed to write solutions formulated by codes based on the example outputs, which is one common scenario in the real-world working process, using examples to convey knowledge.
@@
To obtain large-scale and diverse Case2Code data, we first gather a diverse collection of executable code texts that cover a wide range of real-world applications. Then, we generate the input-output transformation cases with the assistance of LLMs and code interpreters, which do not require powerful LLMs with advanced reasoning capabilities, resulting in a strong-to-weak distilling process. By incorporating LLMs to write input examples for each program and execute the program with these inputs to gather the corresponding outputs, we can synthesize large-scale Case2Code samples with diverse data transformations and complicated control logic.
@@
Based on the synthetic data, we can form a unique and challenging task to evaluate and further train the LLMs and study the induction reasoning ability of LLMs. In the Case2Code challenge, we first test how current LLMs perform in making inductive reasoning. We then train LLMs with Case2Code data to further study whether such data can improve the induction reasoning ability and generalize to other commonly used reasoning tasks. Experimental results show that Case2Code is a challenging task for LLMs, even for powerful LLMs like LLaMA3-70B, GPT-3.5, and GPT-4. With constructed Case2Code data, we can boost LLMs to learn to make such inductive reasoning, while such ability can be transferred to help improve general reasoning tasks such as HumanEval and MBPP in code generation.
@@
(1) We introduce an induction reasoning task for LLMs, Case2Code, pointing out the necessity to synthesize inductive reasoning data for LLMs.
(2) We can teach LLMs to make induction reasoning, improving open-source LLMs by a great margin in the Case2Code challenge.
(3) Equipped with our proposed Case2Code ability, open-source LLMs can be further improved in general reasoning tasks.
@@ 2 Related Work
2.1 Inductive Reasoning
While Reasoning is one major topic for neural networks, especially in the era of LLMs Huang and Chang (2022), inductive reasoning is rarely discussed in LLM reasoning, most research focuses on specific scenarios with limited inductive reasoning. One pioneer work is prerequisite toy tasks Weston et al. (2015) where the task goal is to solve simple induction. Later, Yang et al. (2022) introduces various world-wide knowledge such as botany, history and geography into the facts given and asks neural models to predict whether a given rule is correct. On the other hand, several works focus on training inductive program synthesis models for constrained scenarios with limited search spaces, such as operations on list, string, and manually-defined objects Balog et al. (2016); Devlin et al. (2017); Ellis et al. (2021); Shi et al. (2023). Different from previous reasoning ability studies, our proposed Case2Code task leverages diverse code in the real world as a powerful platform for LLMs to learn inductive reasoning under various challenging scenarios.
2.2 Synthetic Reasoning Data
The most widely studied LLMs reasoning task is the deductive reasoning ability, represented by chain-of-thoughts (CoTs) reasoning Wei et al. (2022b), which instructs LLMs to solve problems in detailed deductive steps. Recent works focus on building high-quality CoTs through strong LLMs such as GPT-4 to enhance smaller LLMs Yu et al. (2023); Mitra et al. (2024); Luo et al. (2023). While a particular line of work focuses on studying different search strategies of reasoning paths, including self-consistency Wang et al. (2022a), rejection sampling Huang et al. (2023); Yuan et al. (2023); Wang et al. (2023a), tree-structure CoT (ToT) searching Yao et al. (2023), Monte Carlo Tree Searching Silver et al. (2016); Chen et al. (2024), etc.
@@ 4.1 Experimental Setup
4.2 Zero-shot Case2Code Performance
@@
4.3 Generalization of Case2Code
@@ 4.4 Ablation Study
Prompt Diversity
Since the synthetic Case2Code training data is converted by triples of (programs, inputs, outputs), during the construction, the prompt templates are utilized to embed the input-output pairs to form natural language texts for LLM to learn. As the LLM can only rely on these converted prompts to learn the Case2Code, it is important to understand the effectiveness of how different prompt templates affect the training of LLMs. Intuitively, the diversity of prompt templates plays an important role in the learning of LLMs. Therefore, we compare synthetic data prompted using a single template style with data utilizing diverse styles of templates. The result is reported in Figure 3, in which diverse prompts may have little effect on the in-domain Case2Code performance, however, the diversity significantly affects the accuracy of LLMs on out-of-domain program synthesis tasks. It is indicated that diversity can be critical during LLM learning, which also has been discussed in other domains like in general natural language processing tasks Wei et al. (2022a) and alignment Ouyang et al. (2022); Wang et al. (2023b).
@@
LLM for Generating Inputs
@@
Model Scale

### v2

# Case2Code: Scalable Synthetic Data for Code Generation

@@ Abstract
Large Language Models (LLMs) have shown outstanding breakthroughs in code generation. Recent work improves code LLMs by training on synthetic data generated by some powerful LLMs, which can be challenging to scale due to the dependence on a teacher model and high generation costs. In this paper, we focus on synthesizing code data at scale and propose a Case2Code task by exploiting the expressiveness and correctness of programs. Case2Code is an inductive inference task that aims to infer underlying code implementations by observing input-output examples or program behaviors, By incorporating LLMs to generate program inputs, and executing the program with these inputs to obtain the program outputs, we can synthesize diverse and high-quality Case2Code data at scale for training and evaluating code LLMs. Experimental results show that case-to-code induction is challenging for current representative LLMs if they are untrained. Models trained with Case2Code improve performance not only on distribution case-to-code induction but also on various coding-generation tasks, demonstrating the great potential of large-scale synthetic data and inductive learning.111Code and datasets are available at https://github.com/choosewhatulike/case2code.
Case2Code: Scalable Synthetic Data for Code Generation
@@ 1 Introduction
The success of large language models (LLMs), exemplified by GPT-4 OpenAI (2023) has revolutionized the AI community. One of the most impressive abilities of LLMs is the code generation ability, exemplified by writing code blocks or programs to accomplish complex instructions and satisfy user specifications Hong et al. (2024); Zhuo et al. (2024); Sun et al. (2024).
To further improve the performance of code generation with open-source LLMs, recent work adopts synthetic data via self-instruct Wang et al. (2023b); Taori et al. (2023) to distill the strong capability of code generation from teacher models. Specifically, practitioners often collect and devise code instructions and then generate high-quality responses using a powerful teacher LLM to synthesize the training data. Then, the data can be used to fine-tune a weak LLM to bootstrap their code generation ability. For example, Code Alpaca Chaudhary (2023) incorporates ChatGPT to generate 20K code instructions under 21 seed tasks for fine-tuning. WizardCoder Luo et al. (2023) adopts Evol-Instruct to the realm of code, using powerful closed-source LLMs to synthetic high-quality code instruction-following samples to enhance open-sourced LLMs code generation performance.
Despite the success of these code synthetic data, they rely on a powerful teacher for data generation Yu et al. (2023b), which can be challenging to scale up as they are bound by the capability of the teacher and suffer from high costs. Moreover, models can perform poorly on tasks like inductive learning as these instructions are rare in the training corpus (Figure 1). Therefore, in this paper, to synthesize high-quality code data at scale and complement existing code training data, we introduce Case2Code, a diverse and challenging synthetic task for LLMs.
@@
Inspired by inductive inference tasks Balog et al. (2016); Devlin et al. (2017); Ellis et al. (2021), we focus on large-scale data synthesis with programs in the real world. In Case2Code, samples are synthesized from real-world productive functions, which are closer to the actual distribution of general LLM applications and production. Specifically, the Case2Code challenge requires the LLM to infer the underlying program based on several input-to-output cases generated by the real-world program. In Case2Code learning, LLMs are supposed to write solutions formulated by codes based on the example outputs, which is one common scenario in the real-world working process, using examples to convey knowledge.
@@
To obtain large-scale and diverse Case2Code data, we first gather a diverse collection of executable code texts that cover a wide range of real-world applications. Then, we generate the input-output transformation cases with the assistance of LLMs and code interpreters. By incorporating LLMs to write input examples for each program and execute the program with these inputs to gather the corresponding outputs, we can synthesize large-scale Case2Code samples with diverse data transformations and complicated control logic. The data synthetic framework does not require powerful LLMs with advanced code generation capabilities, resulting in the possibility of a weak-to-strong learning paradigm.
@@
Based on the synthetic data, we can form a unique and challenging task to evaluate and further train the LLMs and study the case-to-code induction ability of LLMs. In the Case2Code challenge, we first test how current LLMs perform in making the code induction task. We then train LLMs with Case2Code data to further study whether such data can improve the in-distribution code induction ability and generalize to other commonly used code generation tasks. Experimental results show that Case2Code is a challenging task for LLMs, even for powerful LLMs like LLaMA3-70B, GPT-3.5, and GPT-4. With constructed Case2Code data, we can boost LLMs to learn to make such inductive inference tasks, while such ability can be transferred to help improve general code generation tasks such as HumanEval and MBPP.
@@
(1) We introduce a scalable data synthetic framework that aims to generate high-quality and diverse inductive code generation samples for evaluating and training LLMs, Case2Code.
(2) We evaluate the inductive learning ability of current representative LLMs, demonstrating the necessity of synthesizing inductive data like Case2Code for LLMs.
(3) We explore methods of training LLMs on large-scale Case2Code data, showing not only great improvements on the Case2Code challenge but also a consistent generalization of the trained models in general code generation tasks.
@@ 2 Related Work
2.1 Inductive Inference
Inductive inference is rarely discussed in LLM reasoning, most research focuses on specific scenarios with limited inductive reasoning. One pioneer work is prerequisite toy tasks Weston et al. (2015) where the task goal is to solve simple induction. Later, Yang et al. (2022) introduces various world-wide knowledge such as botany, history, and geography into the facts given and asks neural models to predict whether a given rule is correct. In the realm of code, several works focus on Programming by Example (PBE), which aims to induce a valid program given the expected inputs and the corresponding outputs. These works train and evaluate inductive program synthesis models for constrained scenarios with limited search spaces, such as operations on list, string, and manually-defined objects Balog et al. (2016); Devlin et al. (2017); Ellis et al. (2021); Shi et al. (2023); Wu et al. (2024); Li and Ellis (2024). Different from previous studies, our proposed Case2Code task leverages diverse code in the real world as a powerful platform for LLMs to learn inductive inference under various challenging scenarios.
2.2 Synthetic Data
Recent works focus on building high-quality instruction-following or question-answering training data through strong LLMs such as GPT-4 to enhance smaller LLMs Yu et al. (2023a); Mitra et al. (2024); Luo et al. (2023). While a particular line of work focuses on studying different strategies to diverse the instructions and control the quality of LLM generation, including self-consistency Wang et al. (2022), rejection sampling Huang et al. (2023); Yuan et al. (2023); Wang et al. (2023a), program-of-thought (PoT) Sun et al. (2024), tree-structure CoT (ToT) searching Yao et al. (2023), Monte Carlo Tree Searching Silver et al. (2016); Chen et al. (2024), etc. These methods still require a strong LLM as the teacher with high costs of model inference, limiting the scalability.
@@ 4.1 Experimental Setup
4.2 Zero-shot Case2Code is Challenging for Current LLMs
@@
4.3 Generalization of Training on Case2Code
@@ 4.4 Ablation Study
Prompt Diversity of Training Data
Since the synthetic Case2Code training data is converted by triples of (programs, inputs, outputs), during the construction, the prompt templates are utilized to embed the input-output pairs to form natural language texts for LLM to learn. As the LLM can only rely on these converted prompts to learn the Case2Code, it is important to understand the effectiveness of how different prompt templates affect the training of LLMs. Intuitively, the diversity of prompt templates plays an important role in the learning of LLMs. Therefore, we compare synthetic data prompted using a single template style with data utilizing diverse styles of templates. The result is reported in Figure 3, in which diverse prompts may have little effect on the in-domain Case2Code performance, however, the diversity significantly affects the accuracy of LLMs on out-of-domain program synthesis tasks. It is indicated that diversity can be critical during LLM learning, which also has been discussed in other domains like in general natural language processing tasks Wei et al. (2022) and alignment Ouyang et al. (2022); Wang et al. (2023b).
| HumanEval | HumanEval+ | Case2Code | |
| LLaMA3-8B | 35.4 | 20.1 | 29.2 |
| w/ Case2Code (full) | 47.6 | 40.9 | 42.6 |
| w/ Case2Code (code only) | 38.41 | 28.66 | 28.2 |
@@
Importance of I/O pairs
We extract code data from the case2code training data and remove other parts like prompts form the training set, Case2Code (code only). Then, we train Llama3-8B under the same setups of the mixed pre-training in Section 4.3.2, replacing the Case2Code (full) training set with the Case2Code (code only) set. Compared with mixed pre-trained models with Case2Code (code only) dataset, the Case2Code (full) trained models achieve much higher accuracy both on in-distribution Case2Code task and out-of-distribution tasks like HumanEval and HumanEval+, demonstrating the effectiveness of combining the I/O messages with the corresponding code and the incorporation of synthetic case2code samples.
Choice of Input Generator
@@
Improvements on Different Model Scale
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62236004).